# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v3

@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT. In panel (f), the dashed lines represent the situation when the total FR loss is ablated. More ablation studies of visual encoder architecture and handcrafted text prompt can be referred to in Appendix D.
@@
Number of shots per class. From Figure 3(a), we surprisingly find that only 1 shot can make the ASR of Blended and BadNet attacks nearly 0 and significantly decrease the ASR of BadCLIP and increasing the number of shots further enhances the defense. For BadCLIP, 8 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving.
@@
Balancing factor $\alpha$. We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3(b), we observe that both ASR and CA decrease as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of in-dataset predictive features (the adversarial environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small $\alpha$ is sufficient to achieve very low ASR for every evaluated dataset.
@@
Visual context length. We show RVPT’s performance in Figure 3(c) with changing visual context length $b$. It indicates that both ASR and CA are optimal at a context length of approximately 25 to 50 tokens. Beyond this optimal range, CA decreases due to overfitting caused by the increasing number of parameters. Moreover, when the length is 5, the ASRs of all attacks drop below 5%, indicating that tuning only 0.5‱ of parameters is sufficient to defend against all attacks, demonstrating the efficiency of our method. However, with increasing length, the defensive ability is bounded.
@@
Depth of the visual prompt. Figure 3(d) shows the performance as the depth of the features into which the visual prompts are embedded is varied. It is important to note that “depth” here refers to the number of layers counted from the model’s output feature. We also conduct experiments where prompts are stacked from the input feature, as shown in Figure 3(e). The comparison demonstrates that the later layers are more significant for backdoor robustness and recognition performance. We conjecture the reason is that the non-essential features encoded in the earlier layers tend to be sparse and intermixed with desirable features and repelling these features at an earlier stage is difficult and may inadvertently compromise the encoding of discriminative features, resulting in worse performance. Therefore, unless otherwise specified, we will count the depth from the output feature. Moreover, the results in Figure 3(d) indicate the improving performance as the number of the prompted layers increases. Particularly, when the depth reaches approximately six layers, both CA and ASR attain their optimal values.
@@
Margin of the repelling strength. To evaluate the potential risks of over-repelling, we set a margin $m$ of FR Loss $\mathcal{L}_{\mathrm{FRM}_{i}}$ for each sample $\boldsymbol{x}_{i}$: $\mathcal{L}_{\mathrm{FRM}_{i}}=\mathrm{max}(\mathcal{L}_{\mathrm{FR}_{i}},m)$. A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion (which is VPT). The performance of RVPT of different $m$ is shown in Figure 3(f). We observe that both ASR and CA will increase with larger $m$. Therefore, for the purpose of achieving superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
@@ 4.5 Further Explorations
Attention map and t-SNE. We analyze the attention maps and t-SNE JMLR:v9:vandermaaten08a plots of different models. In Figure 5, RVPT enables the encoder to focus on relevant bird features for clean images, while the backdoored model attends to irrelevant ones. For poisoned images, RVPT notices the trigger but avoids overemphasizing it, unlike the backdoored encoder. Figure 5 shows that backdoored CLIP and VPT models cluster around the trigger pattern, ignoring original image content bansal2023cleanclip . RVPT, however, disrupts this cluster and restores poisoned image representations, reducing the influence of the trigger.
RVPT endows CLIP with backdoor robustness in various downstream tasks. To further evaluate the impact of backdoored CLIP in real-world settings, we assess its performance on two downstream tasks: image captioning and text-to-image retrieval in Table 7. For image captioning, we use MSCOCO lin2014microsoft and LLaVA-1.5 (13B) liu2023llava equipped with backdoored CLIP ViT/L14. A successful attack is defined as generating a caption that includes the target class. For text-to-image retrieval, we use an enriched ImageNet-1K deng2009imagenet ; huggingface2025vl with backdoored CLIP (ViT/B32). We insert one poisoned image into the candidate set and if it appears in the top-10 matches for clean captions of the target class, we count it as a successful attack. Table 7 shows that backdooring CLIP during pre-training can transfer the backdoor behavior to the evaluated downstream tasks. Notably, our RVPT effectively mitigates backdoor effects while preserving clean performance, outperforming baselines bansal2023cleanclip ; jia2022visual .
[…]
@@
[…]
@@ 5 Conclusion
Appendix
We summarize the Appendix as follows:
  • Appendix A provides more details of the experimented perturbations and results of PR in CLIP backdoored by more attacks.
  • Appendix B provides more detail in the experiment’s settings of backdoor defense and attack methods.
    • Section B.1 provide the detailed settings of backdoor attacks.
    • Section B.2 provide the detailed settings of backdoor defense baselines.
    • Section B.3 provide the detailed settings of adopted datasets.
  • Appendix C compares computational expenses between our method and CleanCLIP.
  • Appendix D contains more ablation studies, such as the visual encoder architecture, and the handcrafted prompt. It also contains the setting of experiments that ablates the FR loss.
    • Section D.1 provide the ablation study for handcrafted prompt.
    • Section D.2 provide the ablation study for visual encoder architecture.
  • Appendix F discusses both the positive and negative of potential impact of our RVPT.
  • Appendix G discusses limitations of the proposed RVPT.
Appendix A More Results of Perturbation Resistivity of More Backdoored CLIP
  • Gaussian Noise with a standard deviation of 0.001.
  • Gaussian Blur with a kernel size of 3 and variance of 5.
  • Color Jitter with brightness of 0.5 and hue of 0.3
  • Random Horizontal Flip.
The visualizations of these perturbations and trigger patterns can be found in Figure 6 and 9.
We evaluate the perturbation resistivity of four models:
  • ViT/B32 of OpenAI CLIP.
  • ViT/B32 of OpenAI CLIP tuned by VPT with 16 shots on ImageNet.
  • ViT/B32 of OpenAI CLIP tuned by RVPT with 16 shots on ImageNet.
  • ViT/B32 supervised exclusively on ImageNet.
The visual illustration for backdoored CLIP by BadNet and Blended are shown in Figure 7, 8 respectively. The results show that CLIP, even after VPT has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios.
Figure 7: Average PR values of CLIP backdoored by BadNet.
Figure 8: Average PR values of CLIP backdoored by Blended.
Appendix B Detailed Settings
B.1 Detailed Settings of Backdoor Attacks
  • TrojVQA proposes a dual-key backdoor attack that sets triggers on both modalities. To evaluate RVPT under backdoored generative VLMs, we use TrojVQA as the backdoor attack and use the following settings. Firstly, we optimize the visual triggers for the CLIP visual backbone with the description “This is a yellow banana.”. Specifically, We set the patch size to 16$\times$16 and located the patch in the middle of the image. Then, we randomly corrupt the text for poisoned samples into sentences that describe the target class of banana. Moreover, we add the trigger word “SUDO ” in front of each sentence to make sure two keys for one backdoor. In the inference stage, the optimized trigger is added to the image, and “SUDO ” is added to the beginning of the prompt.
We implement the attacks following [3]. Specifically, we employ the AdamW optimizer with an initial learning rate of 1e-6 for all the backdoor attack methods. The learning rate follows a cosine decay schedule over five epochs, and we set the batch size to 128. The visualization of each trigger pattern is in Figure 9.
B.2 Detailed Settings of Backdoor Defenses
B.3 Detailed Settings of Datasets
Appendix C Computational Expenses
We conducted the main experiments on eight single NVIDIA 3090 GPUs using a half-precision data type and recorded one training process in Table 8. The results indicate that our method substantially reduces computational costs, enhancing defense efficiency.
Method Training time GPU memory used Tunable parameters Training samples
CleanCLIP 3:53:02 17640 MB 126 M 250K
RVPT on ImageNet 45:05 3382 MB 0.34 M 16 K
RVPT on OxfordPets 2:19 1010 MB 0.34 M 0.6 K
Appendix D More Experiments for Ablation Study
D.1 Ablation Study of the Handcrafted Prompt
We experiment with the handcrafted prompt of “<CLS>” and “###### <CLS>” and show the result in Table 9. From the result, we can conclude that RVPT still successfully defends against backdoor attacks with different handcrafted prompts.
[…]
       Handcrafted Prompt        ImageNet
       BadNet        blended        BadCLIP
       “<CLS>”        0.09 (62.28)        0.03 (62.20)        3.72 (61.73)
       “###### <CLS>”        0.05 (62.18)        0.02 (61.96)        3.78 (61.37)
       “a photo of a <CLS>”        0.05 (62.76)        0.02 (62.36)        2.76 (61.81)
D.2 Ablation Study of the Visual Encoder Architecture
To evaluate RVPT across different architectures, We poisoned different architectures of CLIP using the same hyper-parameters and show the result in Table 10. In particular, for the sake of fairness, we keep the RVPT training setting for ViT-B/16 and ViT-L/14 the same as for ViT-B/32. The result shows that all attacks have been launched successfully, so we can conclude that RVPT still successfully defends against backdoor attacks in various visual encoder architectures.
[…]
       Method        BadNet        Blended        BadCLIP
       ViT-B/16
       No defense        99.61 (68.04)        98.30 (67.84)        99.67 (68.09)
       Linear Probe        7.11 (67.52)        0.21 (67.30)        97.73 (67.60)
       RVPT        0.66 (68.51)        0.02 (68.32)        3.38 (68.42)
       ViT-L/14
       No defense        98.76 (73.88)        99.02 (74.25)        99.87 (74.68)
       Linear Probe        23.48 (72.31)        0.65 (72.35)        99.35 (72.35)
       RVPT        2.08 (73.60)        0.05 (73.26)        0.59 (74.04)
Appendix E Discussion about the Difference between RVPT and Adversarial Training
As illustrated in Section 3, RVPT adopts the FR loss to adversarially repel features to impede the learning process, which partly shares some similarity with adversarial training (AT) [2]. While this approach shares some superficial similarities with adversarial training (AT) [2], there are fundamental differences between the two methodologies.
As a result, AT prevents the model from encoding the predictive, brittle features, while RVPT encourages the model to encode the in-dataset predictive features. In the training process of RVPT, there are no adversarial features unlearned. Therefore, it cannot defend against adversarial attacks. Table 11 compares the performance of RVPT and a representative adversarial training method of CLIP under adversarial attack. However, there are non-predictive features or out-of-dataset features unlearned, which ensures RVPT to defend against backdoor attacks.
No defense RVPT Adversarial Prompt Tuning
Accuracy 4.37 4.10 13.97
Appendix F Broader Impacts
Positive Impact
Negative Impact
Appendix G Limitations

### v4

@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT. In panel (f), the dashed lines represent the situation when the total FR loss is ablated. More ablation studies of visual encoder architecture and handcrafted text prompt can be referred to in Appendix˜E.
@@
Number of shots per class. From Figure˜3(a), we surprisingly find that only 1 shot can make the ASR of Blended and BadNet attacks nearly 0 and significantly decrease the ASR of BadCLIP and increasing the number of shots further enhances the defense. For BadCLIP, 8 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving.
@@
Balancing factor $\alpha$. We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure˜3(b), we observe that both ASR and CA decrease as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of in-dataset predictive features (the adversarial environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small $\alpha$ is sufficient to achieve very low ASR for every evaluated dataset.
@@
Visual context length. We show RVPT’s performance in Figure˜3(c) with changing visual context length $b$. It indicates that both ASR and CA are optimal at a context length of approximately 25 to 50 tokens. Beyond this optimal range, CA decreases due to overfitting caused by the increasing number of parameters. Moreover, when the length is 5, the ASRs of all attacks drop below 5%, indicating that tuning only 0.5‱ of parameters is sufficient to defend against all attacks, demonstrating the efficiency of our method. However, with increasing length, the defensive ability is bounded.
@@
Depth of the visual prompt. Figure˜3(d) shows the performance as the depth of the features into which the visual prompts are embedded is varied. It is important to note that “depth” here refers to the number of layers counted from the model’s output feature. We also conduct experiments where prompts are stacked from the input feature, as shown in Figure˜3(e). The comparison demonstrates that the later layers are more significant for backdoor robustness and recognition performance. We conjecture the reason is that the non-essential features encoded in the earlier layers tend to be sparse and intermixed with desirable features and repelling these features at an earlier stage is difficult and may inadvertently compromise the encoding of discriminative features, resulting in worse performance. Therefore, unless otherwise specified, we will count the depth from the output feature. Moreover, the results in Figure˜3(d) indicate the improving performance as the number of the prompted layers increases. Particularly, when the depth reaches approximately six layers, both CA and ASR attain their optimal values.
@@
Margin of the repelling strength. To evaluate the potential risks of over-repelling, we set a margin $m$ of FR Loss $\mathcal{L}_{\mathrm{FRM}_{i}}$ for each sample $\boldsymbol{x}_{i}$: $\mathcal{L}_{\mathrm{FRM}_{i}}=\mathrm{max}(\mathcal{L}_{\mathrm{FR}_{i}},m)$. A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion (which is VPT). The performance of RVPT of different $m$ is shown in Figure˜3(f). We observe that both ASR and CA will increase with larger $m$. Therefore, for the purpose of achieving superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
@@ 4.5 Further Explorations
Attention map and t-SNE. We analyze the attention maps and t-SNE JMLR:v9:vandermaaten08a plots of different models. In Figure˜5, RVPT enables the encoder to focus on relevant bird features for clean images, while the backdoored model attends to irrelevant ones. For poisoned images, RVPT notices the trigger but avoids overemphasizing it, unlike the backdoored encoder. Figure˜5 shows that backdoored CLIP and VPT models cluster around the trigger pattern, ignoring original image content bansal2023cleanclip . RVPT, however, disrupts this cluster and restores poisoned image representations, reducing the influence of the trigger.
RVPT endows CLIP with backdoor robustness in various downstream tasks. To further evaluate the impact of backdoored CLIP in real-world settings, we assess its performance on two downstream tasks: image captioning and text-to-image retrieval in Table˜7. For image captioning, we use MSCOCO lin2014microsoft and LLaVA-1.5 (13B) liu2023llava equipped with backdoored CLIP ViT/L14. A successful attack is defined as generating a caption that includes the target class. For text-to-image retrieval, we use an enriched ImageNet-1K deng2009imagenet ; huggingface2025vl with backdoored CLIP (ViT/B32). We insert one poisoned image into the candidate set and if it appears in the top-10 matches for clean captions of the target class, we count it as a successful attack. Table˜7 shows that backdooring CLIP during pre-training can transfer the backdoor behavior to the evaluated downstream tasks. Notably, our RVPT effectively mitigates backdoor effects while preserving clean performance, outperforming baselines bansal2023cleanclip ; jia2022visual .
Table 7: Results on image captioning (IC) and text-to-image retrieval (TR). For IC, we evaluate backdoor robustness using ASR ($\downarrow\%$), defined as the percentage of generated captions that include the target class. Clean performance is measured using the BLEU score ($\uparrow\%$). For TR, ASR ($\downarrow\%$) is defined as the percentage of target-class captions where the poisoned image appears in the top-10 most similar images to evaluate the backdoor robustness. We also report the average rank ratio ($\downarrow\%$) of the poisoned image among target-class candidates. Clean accuracy ($\uparrow\%$) refers to the percentage of original captions correctly matched with their corresponding images (top-1 similarity).
@@
Attack ASR ($\downarrow\%$) BLEU ASR ($\downarrow\%$) avg. Rank Ratio ($\uparrow\%$) CA ($\%$)
@@ 5 Conclusion
Acknowledgment
This work is supported by the Big Data Computing Center of Southeast University. Haobo Wang is supported by the Fundamental Research Funds for the Central Universities (No. 226-2025-00004).
@@ 80. Guidelines:
Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning
Appendix
We summarize the Appendix as follows:
  • Appendix˜A provides a detailed reason of using VPT to defend backdoored CLIP.
  • Appendix˜B provides more details of the experimented perturbations and results of PR in CLIP backdoored by more attacks.
  • Appendix˜C provides more detail in the experiment’s settings of backdoor defense and attack methods.
    • Section˜C.1 provide the detailed settings of backdoor attacks.
    • Section˜C.2 provide the detailed settings of backdoor defense baselines.
    • Section˜C.3 provide the detailed settings of adopted datasets.
  • Appendix˜D compares computational expenses between our method and CleanCLIP.
  • Appendix˜E contains more ablation studies, such as the visual encoder architecture, and the handcrafted prompt. It also contains the setting of experiments that ablates the FR loss.
    • Section˜E.1 provide the ablation study for handcrafted prompt.
    • Section˜E.2 provide the ablation study for visual encoder architecture.
  • Appendix˜G discusses both the positive and negative of potential impact of our RVPT.
  • Appendix˜H discusses limitations of the proposed RVPT.
Appendix A Why Using VPT to defend backdoored CLIP?
A.1 The drawback of fine-tuning the whole model
Fully fine-tuning can be done in two ways, both of which exhibit significant drawbacks in our setting.
Fine-tuning whole CLIP with Info-NCE loss. Since our primary downstream task is classification, we do not have downstream image-text pairs for optimization. We could use surrogate datasets such as CC3M, but as Table 8 shows, defense performance drops significantly. This degradation occurs because RVPT relies on identifying predictive features specific to the downstream dataset and pruning non-predictive ones. The mismatch between predictive features in CC3M and the actual downstream task leads to poor backdoor robustness.
Fine-tuning whole CLIP with cross-entropy loss. This requires adding an additional linear layer on top of the visual encoder for closed-set classification. However, this approach has several drawbacks: (i) It removes CLIP’s open-set classification capability, severely limiting its generalization to other datasets. (ii) Given the limited labeled data, the large number of tunable parameters will cause overfitting, often resulting in significant catastrophic forgetting.
In summary, fine-tuning the whole CLIP with InfoNCE loss yields poor backdoor robustness and higher computational cost, while using cross-entropy loss slightly improves robustness but at the expense of much greater data requirements and loss of model generalization. Neither option provides a favorable trade-off between robustness, efficiency, and generalization.
Table 8: ASR (CA) performance of different fine-tuning methods with or without FR loss.
Fine-tuning method without FR Loss with FR Loss Open-set capability
ImageNet with CE Loss 10.09 (60.71) 0 (59.22) NO
CC3M with Info-NCE Loss 92.50 (54.50) 18.23 (54.66) YES
VPT on few-shot ImageNet (preferred) 98.14 (64.32) 2.76 (61.81) YES
A.2 The reasons why we use VPT
As outlined above, full fine-tuning of CLIP presents significant drawbacks, either in terms of degraded robustness, excessive data requirements, or loss of open-set capabilities. To address these challenges, we adopt parameter-efficient fine-tuning methods (VPT) that:
  • Leverage limited labeled downstream data effectively, aligning with RVPT’s core mechanism of identifying task-relevant visual features.
  • Preserve CLIP’s open-set classification ability, which is critical for generalization across diverse tasks.
Specifically, since our goal is to enhance the perturbation resistivity of the CLIP’s visual features, we inject learnable parameters into the visual modality, leading to our choice of VPT. This approach significantly improves downstream backdoor robustness using much less data while retaining CLIP’s open-set capability, as demonstrated in Table 8.
Appendix B More Results of Perturbation Resistivity of More Backdoored CLIP
  • Gaussian Noise with a standard deviation of 0.001.
  • Gaussian Blur with a kernel size of 3 and variance of 5.
  • Color Jitter with brightness of 0.5 and hue of 0.3
  • Random Horizontal Flip.
The visualizations of these perturbations and trigger patterns can be found in Figure˜6 and 9.
We evaluate the perturbation resistivity of four models:
  • ViT/B32 of OpenAI CLIP.
  • ViT/B32 of OpenAI CLIP tuned by VPT with 16 shots on ImageNet.
  • ViT/B32 of OpenAI CLIP tuned by RVPT with 16 shots on ImageNet.
  • ViT/B32 supervised exclusively on ImageNet.
The visual illustration for backdoored CLIP by BadNet and Blended are shown in Figure˜7, 8 respectively. The results show that CLIP, even after VPT has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios.
Figure 7: Average PR values of CLIP backdoored by BadNet.
Figure 8: Average PR values of CLIP backdoored by Blended.
Appendix C Detailed Settings
C.1 Detailed Settings of Backdoor Attacks
  • TrojVQA proposes a dual-key backdoor attack that sets triggers on both modalities. To evaluate RVPT under backdoored generative VLMs, we use TrojVQA as the backdoor attack and use the following settings. Firstly, we optimize the visual triggers for the CLIP visual backbone with the description “This is a yellow banana.”. Specifically, We set the patch size to $\times16$ and located the patch in the middle of the image. Then, we randomly corrupt the text for poisoned samples into sentences that describe the target class of banana. Moreover, we add the trigger word “SUDO ” in front of each sentence to make sure two keys for one backdoor. In the inference stage, the optimized trigger is added to the image, and “SUDO ” is added to the beginning of the prompt.
We implement the attacks following [3]. Specifically, we employ the AdamW optimizer with an initial learning rate of 1e-6 for all the backdoor attack methods. The learning rate follows a cosine decay schedule over five epochs, and we set the batch size to 128. The visualization of each trigger pattern is in Figure˜9.
C.2 Detailed Settings of Backdoor Defenses
C.3 Detailed Settings of Datasets
Appendix D Computational Expenses
We conducted the main experiments on eight single NVIDIA 3090 GPUs using a half-precision data type and recorded one training process in Table˜9. The results indicate that our method substantially reduces computational costs, enhancing defense efficiency.
Method Training time GPU memory Parameters Training samples
CleanCLIP 3:53:02 17640 MB 126 M 250K
RVPT on ImageNet 45:05 3382 MB 0.34 M 16 K
RVPT on OxfordPets 2:19 1010 MB 0.34 M 0.6 K
Appendix E More Experiments for Ablation Study
E.1 Ablation Study of the Handcrafted Prompt
We experiment with the handcrafted prompt of “<CLS>” and “###### <CLS>” and show the result in Table˜10. From the result, we can conclude that RVPT still successfully defends against backdoor attacks with different handcrafted prompts.
Table 10: Performance of RVPT with different handcrafted prompts. We report ASR ($\downarrow\%$), with CA ($\uparrow\%$) shown in parentheses. We can see the RVPT is stable regarding different handcrafted prompts.
   Handcrafted Prompt    ImageNet
   BadNet    blended    BadCLIP
   “<CLS>”    0.09 (62.28)    0.03 (62.20)    3.72 (61.73)
   “###### <CLS>”    0.05 (62.18)    0.02 (61.96)    3.78 (61.37)
   “a photo of a <CLS>”    0.05 (62.76)    0.02 (62.36)    2.76 (61.81)
E.2 Ablation Study of the Visual Encoder Architecture
To evaluate RVPT across different architectures, We poisoned different architectures of CLIP using the same hyper-parameters and show the result in Table˜11. In particular, for the sake of fairness, we keep the RVPT training setting for ViT-B/16 and ViT-L/14 the same as for ViT-B/32. The result shows that all attacks have been launched successfully, so we can conclude that RVPT still successfully defends against backdoor attacks in various visual encoder architectures.
Table 11: Performance of RVPT on attacked samples of ImageNet across different visual architectures. We report ASR ($\downarrow\%$), with CA ($\uparrow\%$) shown in parentheses. We can see the RVPT consistently successfully defends against the backdoor attacks in various visual encoder architectures.
   Method    BadNet    Blended    BadCLIP
   ViT-B/16
   No defense    99.61 (68.04)    98.30 (67.84)    99.67 (68.09)
   Linear Probe    7.11 (67.52)    0.21 (67.30)    97.73 (67.60)
   RVPT    0.66 (68.51)    0.02 (68.32)    3.38 (68.42)
   ViT-L/14
   No defense    98.76 (73.88)    99.02 (74.25)    99.87 (74.68)
   Linear Probe    23.48 (72.31)    0.65 (72.35)    99.35 (72.35)
   RVPT    2.08 (73.60)    0.05 (73.26)    0.59 (74.04)
Appendix F Discussion about the Difference between RVPT and Adversarial Training
As illustrated in Section˜3, RVPT adopts the FR loss to adversarially repel features to impede the learning process, which partly shares some similarity with adversarial training (AT) [2]. While this approach shares some superficial similarities with adversarial training (AT) [2], there are fundamental differences between the two methodologies.
As a result, AT prevents the model from encoding the predictive, brittle features, while RVPT encourages the model to encode the in-dataset predictive features. In the training process of RVPT, there are no adversarial features unlearned. Therefore, it cannot defend against adversarial attacks. Table˜12 compares the performance of RVPT and a representative adversarial training method of CLIP under adversarial attack. However, there are non-predictive features or out-of-dataset features unlearned, which ensures RVPT to defend against backdoor attacks.
No defense RVPT Adversarial Prompt Tuning
Accuracy 4.37 4.10 13.97
Appendix G Broader Impacts
Positive Impact
Negative Impact
Appendix H Limitations